OpenStudy (anonymous):

Solve the following systems of equations using the addition (elimination) method. What type of system is it? Name the solution if there is one. 5x+y=2 3x+y=1

OpenStudy (anonymous):

In order to solve the system of equations make sure that when you add or subtract them one of the variable is removed that is when subtracted or added because of equal coefficient they get cancelled and other is solvable by using the constant, this is addition elimination to the best of my knowledge

OpenStudy (anonymous):

In you question just subtract them , resultant is (5x+y)-(3x+y)=2-1 => 2x=1 => x = 0.5 (1/2) and y = -0.5

OpenStudy (anonymous):

so then it would be inconsistent?

OpenStudy (anonymous):

@ritesh_asu

OpenStudy (anonymous):

no it is not inconsistent, we have solved and got the answer right ?

OpenStudy (anonymous):

In inconsistent we won't be able to get an answer, something like x+2y=2 and 2x+4y=4

OpenStudy (anonymous):

okay so since we did get an answer of x=.5 and y=-.5 it is consistent

OpenStudy (anonymous):

@ritesh_asu

OpenStudy (anonymous):

yes , check this reading , it might help you http://www.wyzant.com/resources/lessons/math/precalculus/systems_of_equations/consistent_and_inconsistent_systems

OpenStudy (anonymous):

@spencerbeymer

OpenStudy (anonymous):

awesome thanks @ritesh_asu